The sum of an interger and twice the next consecutive integer is 29. What are the integers?
Answer by ptaylor(2198)

Let x=the integer
Then x+1=the next consecutive integer
2(x+1)=twice the next consecutive integer
Now we are told that:
x+2(x+1)=29 get rid of parens
x+2x+2=29 subtract 2 from both sides
x+2x+2-2=29-2 collect like terms
3x=27 divide both sides by 3
x=9-----------------------the integer
x+1=9+1=10------------------next consecutive integer
CK
9+2*10=29
29=29
Hope this helps-----ptaylor